The number of work-related injuries per month in a manufacturing plant is known to follow a Poisson distribution, with a mean of 2.5 work-related injuries a month. What is the probability that in a given month, no work-related injuries occur? That at least one work-related injury occurs?


1
Expert's answer
2022-04-12T06:44:27-0400

We have a Poisson distribution,

"\lambda=2.5;\\
P_t(X=k)=\cfrac{(\lambda t)^k\cdot e^{-\lambda t}}{k!}."


"1) P_{1}(X=0)=\cfrac{(2.5\cdot1)^0\cdot e^{-2.7\cdot1}}{0!}=0.0821;\\
2) P_{1}(X\ge1)=1-P(X<1)=1-P(X=0)=\\
=1-0.0821=0.9179."